OpenStudy (wasiqss):

solve this Differential equation by finding Integrating factor (5xy+4y^2+1)dx+(x^2+2xy)dy=0

OpenStudy (wasiqss):

ash start

OpenStudy (ash2326):

Wasiqss I'm thinking about it, wait for some time:)

OpenStudy (ash2326):

I think I'm gonna get disconnected soon, but I'll solve it for you. We'll have power cut now:(

OpenStudy (wasiqss):

solve it man plz

OpenStudy (wasiqss):

lalaly any idea? :P

OpenStudy (lalaly):

let M=5xy +4y^2+1 N=x^2+2xy \[M_y=5x+8y\]\[N_x=2x+2y\] now we ned to find the integratiing factor\[\frac{1}{F} \frac{dF}{dx}=\frac{M_y-N_x}{N}\]\[\frac{1}{F} \frac{dF}{dx}=\frac{5x+8y-(2x+2y)}{x^2+2xy}\]\[\frac{1}{F} \frac{dF}{dx}=\frac{3(x+2y)}{x(x+2y)}\]\[\frac{1}{F} \frac{dF}{dx}=\frac{3}{x}\]integrate to find F which is the integrating factor Then multiply the whole equation by F and then its an exact differential equation do u know how to find the solution of an exact diff equation?

OpenStudy (wasiqss):

lalaly thanku thanku :) sister u rock :D

OpenStudy (lalaly):

lol anytime:)